# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v7

@@ A.3.2 Geometric Properties of Global Minima
In our specific case, since $f$ has a global minimum at 0, $w$* also represents the lowest point of the epigraph along the $(n+1)$st axis. Our proofs utilize the following result:
@@
The above relationship connects naturally to conjugate function theory. As shown in Section 4.2.1 of Bertsekas (2009), for the MC/MC framework we have $w^{*}=f(0)$ and $q^{*}=f^{**}(0)$, providing another perspective on our result: for closed proper convex functions, the equality $q^{*}=w^{*}$ aligns with the fundamental property that $f(x)=f^{**}(x)$ for all $x\in\mathbb{R}^{n}$.
These properties, established for minimum at 0, extend naturally to any point $c\in\mathbb{R}^{n}$ through translation. Since translation moves points parallel to coordinate axes without rotation, it preserves all geometric relationships between hyperplanes and the epigraph, particularly the non-verticality property. We formalize this in the following proposition:
@@
Let $f$ be a convex function with finite global minimum at $c$. Then $q^{*}=w^{*}=f(c)$.
@@
Slater’s condition follows trivially in our setup as shown in the proposition below
@@ A.3.4 A plausible explanation for the non-convergence of the ReLU networks with the Capture Theorem
which is independent of the Lagrange multiplier, which implies that there is no dual "force" anchoring the iterates towards the optimal solution. We hypothesize that this is the reason why our NDGM sequence in our Capture Theorem plots are bouncing between "equally plausible" points. To use the "taxation" analogy of the dual variables (section 5.4.4 in Boyd et al. (2004)), the price of violation becomes irrelevant when nothing is stored in the warehouse!

### v8

@@ A.3.2 Geometric Properties of Global Minima
In our specific case, since $f$ has a global minimum at 0, $w^{*}$ represents the global lowest point on the epigraph an represents the lowest point of the epigraph along the $(n+1)$st axis. Our proofs utilize the following result:
@@
The above relationship connects naturally to conjugate function theory. As shown in Section 4.2.1 of Bertsekas (2009), for the MC/MC framework we have $w^{*}=f(0)$ and $q^{*}=f^{**}(0)$, providing another perspective on our result: for closed proper convex functions, the equality $q^{*}=w^{*}$ aligns with the fundamental property that $f(x)=f^{**}(x)$ for all $x\in\mathbb{R}^{n}$. By enforcing some mild constraints on the nature of $f$, we can provide an even stronger qualification of the dual optimal solution as is shown by the proposition below
@@
Let $f$ be a convex function with a finite global minimum at 00. Then the set of solutions for the max crossing problem are non empty and compact if and only if $f$ is continuous around 0.
Proof.
First suppose $f$ is continuous at 0. Non emptiness is established in the previous propositions. To establish compactness (Propositions 4.4.1 and 4.4.2 in Bertsekas (2009)) the following three conditions need to be satisfied.
  • $w^{*}>-\infty$
  • $\overline{M}=M+\left\{(0,w)|w\geq 0\right\}$ is convex
  • The origin is an interior point of the set $D=\left\{u\text{ | there exists }w\in R\text{ with }(u,w)\in\overline{M}\right\}$
The first condition follows immediately since $f$ has a finite minimum at 0. The second condition holds because $M=\text{Epi}(f)$ is convex (as $f$ is convex), and for any $(x,w)\in M$ and $e\geq 0$, we have $(x,w+e)\in M$, so $\overline{M}=M$.
For the third condition, since $f$ is continuous around the origin, it will be finite and continuous around the origin (since it is finite at 0). Thus, $D$ contains a ball centered at the origin, and therefore, the origin is an interior point of $D$.
Now, for the reverse direction, suppose that the set of solutions for the max crossing problem are non-empty and compact. Then, from Proposition 4.4.2 of Bertsekas (2009), the origin has to be an interior point of $D$ described in point number 3 above. Since $f$ has a finite global minimum at 0, we know that 1) $f$ is proper, and 2) $0\in dom(f)=\left\{x\in X\text{ | }f(x)<\infty\right\}$. To establish a contradiction, assume that f is discontinuous at 0. This would imply that the origin is not in the relative interior of $dom(f)$ (if $f$ is a proper convex function on (,]($\infty$,$\infty$]( - ∞ , ∞ ], then $f$ restricted to $dom(f)$ is continuous over the relative interior of $dom(f)$ (Proposition 1.3.11 in Bertsekas (2009))). In order for the discontinuity to exist, the function has to “jump” away from 0 in some direction. We consider three possible forms such a jump could take:
  • The highest function jump is to some value value lower than $w^{*}$. This is not possible because $w^{*}$ is the global minimum. To argue this result another way, note that the “highest jump" to a value smaller than $w^{*}$ is not possible because it would violate lower semicontinuity (Proposition A.3)
  • The highest function jump is to some finite value larger than $w^{*}$. This is not possible, because if so, then 0 will become a relative interior point of $dom(f)$, which would require the function to be continuous at 0 thus establishing a contradiction. Mathematically, $0\in dom(f)$, and in this case, there would exist a neighborhood of 00 in every direction in $dom(f)$, and therefore it would be possible to “prolong" 0 within $dom(f)$ without leaving the set. Thus by the prolongation lemma (paraphrasing Proposition 1.3.3 of Bertsekas (2009) to our use case, 0 is a relative interior of a non-empty convex set, $C$, if and only if $0\in C$ and for every $u\in C$, there exists a $\gamma>0$ such that $-\gamma u\in C$), 0 would have to belong to the relative interior of $dom(f)$, which would require the function to be continuous at that point.
  • The highest function jump is $\infty$. This is not possible because it would not be possible to draw a ball around the origin in $dom(f)$, i.e., 0 cannot be an interior point of $D$.
Since all three possibilities lead to contradictions, we conclude that $f$ must be continuous at 0. ∎
The properties established above for minimum at 0, extend naturally to any point $c\in\mathbb{R}^{n}$ through translation. Since translation moves points parallel to coordinate axes without rotation, it preserves all set properties, and the geometric relationships between hyperplanes and the epigraph, particularly the non-verticality property. We formalize this in the following proposition.
Proposition A.7.
Let $f$ be a convex function with a finite global minimum at $c$. Then $q^{*}=w^{*}=f(c)$. Furthermore, if $f$ takes finite values on an open ball around $c$, then the set of optimal solutions of the max crossing problem will be non empty and compact
@@
The following general result follows from the above proposition.
Proposition A.8.
Suppose a continuous, convex function, $f$, has a finite global minimum. Then the solution to the max crossing problem will be non-empty and compact
Proof.
Special case of the previous proposition ∎
As an application of the above proposition, we note that the set of optimal solutions to the max crossing problem for (7) will be non empty and compact. Slater’s condition follows trivially in our setup as shown in the proposition below
@@ A.3.4 A plausible explanation for the non-convergence of the ReLU networks with the Capture Theorem
which is independent of the Lagrange multiplier. There are two conclusions from this: first, the set of dual solutions is unbounded, and is therefore not compact corroborating Proposition A.7 since our modified loss function is not continuous at 0, and second that there is no dual "force" anchoring the iterates towards the optimal solution. This also corroborates with Proposition A.6; since We hypothesize that this is the reason why our NDGM sequence in our Capture Theorem plots are bouncing between "equally plausible" points. To use the "taxation" analogy of the dual variables (section 5.4.4 in Boyd et al. (2004)), the price of violation becomes irrelevant when nothing is stored in the warehouse!